# Tactile Aware Dynamic Obstacle Avoidance in Crowded Environment with Deep Reinforcement Learning

### v1

@@ III-C Reinforcement Learning
Collision query: To introduce a tactile aware intelligence layer, the robot model is equipped with 6 contact sensor plates aimed at detecting collision queries from all directions. The contact sensor input, $[F_{c}]$, is calibrated to correspond with the real-life force sensor readings, with a maximum of 10 kg loading. The reading indicates the amount of load exerted on the robot by contact objects or obstacles.
Obstacle proximity: The laser scan data is a processed output from the package ira_laser_tool which merges the laser scan data from two opposite-facing LiDARs into a full 360° field of view around the robot. The maximum detection range is 30 m. The laser measurements, $[d_{l}]$, are compressed into 18 values via minimum pooling and subjected to Gaussian noise with a standard deviation $\sigma=0.01$.
@@
The reward function awards the agent for reaching the goal and penalizes severely when colliding with dynamic obstacles or static obstacles according to $-K_{f}\times\min(F_{c_{1,2,...,6}})$. In addition, close proximity to dynamic obstacles or objects is penalized slightly to prevent extreme close contact most of the time by $-K_{p}\times(d_{p}-\min(d_{l_{1,2,...,18}}))$ below a threshold distance $d_{p}=0.5\ m$. The robot is subject to a higher penalty if it does not move towards the goal than not facing the goal so as to discourage non-progressive movements. In general, the design of the reward policy is to keep the mean reward range per episode small within [-10,10] where convergence can be expected around 0. The final goal of the agent is to choose actions at each time step in order to maximize its expected future discounted reward, given by:
@@
One training environment is initialized at a time. The RL training is run using the PPO algorithm [30, 26] for 1 million timesteps, where $dt=0.01\ s$ in simulation, with 10 epochs and 32 minibatches. 5 policy evaluation episodes are run every 10,000 training timesteps.
@@ IV-B Simulation Results
[…]
@@
A*-EBand-APF presents good performance in the number of collisions and time with low error in position. The APF is traditionally known to be easily trapped in a local minima especially when more than two obstacles come around the robot agent or when it requires to pass through a tight corridor. The low success rate may be explained by the repulsive force of APF which causes perturbed actions in the proximity of obstacles, leading the robot to deviate largely from its path, which results in the EBand planner experiencing robot pose acquisition problem. Methods like A*-EBand and A*-DWA seems to have a high success rate but experiences especially high volume of collisions because the $move\_base$ planner is not sufficiently reactive, requiring more time to be able to replan especially in dynamic environments with constantly changing costmap, taking a long time to replan its path. A*-EBand-RL on the other hand, overcomes this limitation as the RL velocity trajectory optimizer will not fail to run even when the EBand planner fails in calculating a path and trajectory command, allowing unperturbed navigation.

### v2

@@ III-C Reinforcement Learning
Collision query: To introduce a tactile aware intelligence layer, the robot model is equipped with 6 contact sensor plates aimed at detecting collision queries from all directions. The contact sensor input, $F_{c}]$, is calibrated to correspond with the real-life force sensor readings, with a maximum of 10 kg loading. The reading indicates the amount of load exerted on the robot by contact objects or obstacles.
Obstacle proximity: The laser scan data is a processed output from the package ira_laser_tool which merges the laser scan data from two opposite-facing LiDARs into a full 360° field of view around the robot. The maximum detection range is 30 m. The laser measurements, $d_{l}]$, are compressed into 18 values via minimum pooling and subjected to Gaussian noise with a standard deviation $\sigma=0.01$.
@@
The reward function awards the agent for reaching the goal and penalizes severely when colliding with dynamic obstacles or static obstacles according to $K_{f}\times\min(F_{c_{1,2,...,6}})$. In addition, close proximity to dynamic obstacles or objects is penalized slightly to prevent extreme close contact most of the time by $K_{p}\times(d_{p}-\min(d_{l_{1,2,...,18}}))$ below a threshold distance $d_{p}=0.5\ m$. The robot is subject to a higher penalty if it does not move towards the goal than not facing the goal so as to discourage non-progressive movements. In general, the design of the reward policy is to keep the mean reward range per episode small within [-10,10] where convergence can be expected around 0. The final goal of the agent is to choose actions at each time step in order to maximize its expected future discounted reward, given by:
@@
One training environment is initialized at a time. The RL training is run using the PPO algorithm [30, 26] for 1 million timesteps, where dt=0.01sdt=0.01\ s in simulation, with 10 epochs and 32 minibatches. 5 policy evaluation episodes are run every 10,000 training timesteps.
@@ IV-B Simulation Results
We compare the performance of the different algorithms using the metrics: success rate $\%$], instances of collision, average time taken to reach goal, average path length [$m$] and final position error [$m$] (See Fig. 5). Success rate describes the fraction of cases where the robot agent reaches the goal position without the algorithm failing. The instances of collision is tabulated as 1 in every timestep regardless of whether one or more contact sensors register collision queries in that timestep. Time taken shows how long it took for the robot to reach the goal and average path length indicates how far the robot travelled to reach the goal. Position error is measured as the difference between the robot’s final position and the goal.
@@
A*-EBand-APF presents good performance in the number of collisions and time with low error in position. The APF is traditionally known to be easily trapped in a local minima especially when more than two obstacles come around the robot agent or when it requires to pass through a tight corridor. The low success rate may be explained by the repulsive force of APF which causes perturbed actions in the proximity of obstacles, leading the robot to deviate largely from its path, which results in the EBand planner experiencing robot pose acquisition problem. Methods like A*-EBand and A*-DWA seems to have a high success rate but experiences especially high volume of collisions because the move_basemove\_base planner is not sufficiently reactive, requiring more time to be able to replan especially in dynamic environments with constantly changing costmap, taking a long time to replan its path. A*-EBand-RL on the other hand, overcomes this limitation as the RL velocity trajectory optimizer will not fail to run even when the EBand planner fails in calculating a path and trajectory command, allowing unperturbed navigation.